Lido Company's standard and actual costs per unit for the most recent period, during which 400 units were actually produced, are given below:
Standard Actual
Materials:
Standard: 2 metres at $1.50 per m. $3.00
Actual: 2.1 metres at $1.60 per m. $3.36
Direct Labour:
Standard: 1.5 hours at $6.00 per hr. $9.00
Actual: 1.4 hours at $6.50 per hr. $9.10
Variable overhead:
Standard: 1.5 hours at $3.40 per hr. $5.10
Actual: 1.4 hours at $3.10 per hr. $4.34
Total Unit Cost: $17.10 $16.80
There were no inventory of materials at the beginning or end of the period.
Required:
(a) Materials price variance
(b) Materials quantity variance
(c) Direct labour rate variance
(d) Direct labour efficiency variance
(e) Variable overhead spending variance
(f) Variable overhead efficiency variance
Answer #1
Standard Actual
Particulars Qty/ hours Rate amount Qty/ hours Rate amount
Materials                    800.00                        1.50                1,200.00                    840.00                        1.60                1,344.00
Labour                    600.00                        6.00                3,600.00                    560.00                        6.50                3,640.00
Overhead                    600.00                        3.40                2,040.00                    560.00                        3.10                1,736.00
Actual output                    400.00
DMPV = (SP-AP)*AQ purchased
DMPV = (1.5 - 1.60)840
DMPV = 84 U
DMQV= (SQ-AQ)SP
DMQV= (800 - 840)1.5
DMQV= 60 U
DLRV= (SR-AR)AH
DLRV= (6 - 6.50)560
DLRV= 280 U
DLEV = (SH-AH)SR
DLEV = (600 - 560)6
DLEV = 240 F
VORV = (SR-AR)AH
VORV = (3.4 - 3.1)560
VORV = 168 F
VOEV = (SH-AH)SR
VOEV = (600 -560)3.4
VOEV = 136 F